OpenStudy (anonymous):

How to convert from cubic meter (m^3) to millimetre (mm)?

OpenStudy (mathstudent55):

1 m = 1000 mm (1 m)^3 = (1000 mm)^3

OpenStudy (mathstudent55):

1 m^3 = 10^9 mm^3

OpenStudy (anonymous):

For instance how many mm of rain to fill 30.9 m^3?

OpenStudy (mathstudent55):

You mentioned twice m^3 and mm, not mm^3. To convert m^3 to mm^3, multiply m^3 by 10^9 30.9 m^3 * 10^9 mm^3/m^3 = 3.9 * 10^10 mm^3

OpenStudy (anonymous):

In the text they tell me to convert from m^3 to mm

OpenStudy (anonymous):

Textbook*

OpenStudy (anonymous):

In other words, how many mm of rain to fill three water tanks whose total volume is 30.9m^3?

OpenStudy (anonymous):

It says in the textbook regarding the last question "when rain falls it is measured I'm millimetres. This means that this is the depth to which the water would fill if it were captured. Use the formula V=AH to find the number of millimetres of rain that must fall on the roof to fill the tanks.

OpenStudy (anonymous):

Total volume of the three tanks is 30.9m^3 A rectangular rooftop has dimensions 17m x 10m providing water to the three tanks (each with a radius of 1.25m and a height of 2.1m

OpenStudy (mathstudent55):

What are the dimesions of the roof?

OpenStudy (anonymous):

I sent you the dimensions

OpenStudy (mathstudent55):

Ok, I see it.

OpenStudy (mathstudent55):

You are correct, the total volume of the three tanks is 30.9 m^3

OpenStudy (anonymous):

I solved the first three question and that was one of them

OpenStudy (mathstudent55):

The volume of a prism that has as base the roof has a volume of V = LWH We know the volume is 30.9 m^3, L = 17 m, and W = 10 m; we need to find H

OpenStudy (mathstudent55):

H = V/(LW) H = (30.9 m^3)/(17 m * 10 m)

OpenStudy (mathstudent55):

H = 0.18 m

OpenStudy (anonymous):

30.9m^3/170m^2?

OpenStudy (mathstudent55):

Correct

OpenStudy (anonymous):

I'm with you.

OpenStudy (mathstudent55):

Once you get H, you'll have the height in meters. Then you need to convert it to mm. To do that, keep in mind that 1 m = 1000 mm, so multiply the height in meters by 1000 to get it in mm.

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

Could you please show it to me. I'm not that confident when it comes to converting units

OpenStudy (anonymous):

Also I have one last question that is driving me nuts :)

OpenStudy (mathstudent55):

Did you get H = 0.18 m?

OpenStudy (anonymous):

Yes I understood where it came from thanks to you

OpenStudy (mathstudent55):

Just take the height and multiply it by 1000: 0.18 m * 1000mm/m = 180 mm

OpenStudy (anonymous):

Thanks a lot for your help. I've got a something similar to this question if you don't mind helping me

OpenStudy (mathstudent55):

go ahead. pls start a new post. i'll look for it

OpenStudy (anonymous):

Cheers pal